OpenStudy (anonymous):

The table below shows a direct variation. What is the value of m? x y 12 24 18 m 24 48 2 24 30 36

OpenStudy (trojanpoem):

\[x \alpha y\] x = const y , (let const = k) x = ky when x = 12, y = 24 12 = 24k , k = 0.5 x = 0.5y , y = 2x put x = 18 , y = 2 * 18 = 36

OpenStudy (anonymous):

thanks